### v1

# Integrable open elliptic Toda chain with boundaries

Abstract
In this letter we discuss the classical integrable elliptic Toda chain proposed by I. Krichever. Our goal is to construct an open elliptic Toda chain with boundary terms. This is achieved using the factorized form of the Lax matrix and gauge equivalence with the XYZ chain.
1 Introduction
The elliptic closed Toda chain was introduced by I. Krichever [6]. It is a classical integrable system defined by the following Hamiltonian:
$\begin{array}[]{c}\displaystyle{H^{\hbox{\tiny{eToda}}}=-\frac{1}{2}\sum\limits_{a=1}^{n}\Big(\log\frac{1}{\sinh^{2}({\bf{p}}_{a}/2c)}+\log\Big(\wp({\bf{q}}_{a-1}-{\bf{q}}_{a})-\wp({\bf{q}}_{a-1}+{\bf{q}}_{a})\Big)\Big)\,,}\end{array}$ (1)
where ${\bf{p}}_{a}$, ${\bf{q}}_{a}$, a=1,,na=1,...,n is a set of canonically conjugated variables (that is $\{{\bf{p}}_{a},{\bf{q}}_{b}\}=\delta_{ab}$) and $\wp(x)$ is the Weierstrass elliptic function on elliptic curve with moduli $\tau$, ${\rm Im}(\tau)>0$. We assume here that numeration of particles is modulo $n$: ${\bf{q}}_{n+1}={\bf{q}}_{1}$, ${\bf{q}}_{0}={\bf{q}}_{n}$. This model is a particular case of a more general elliptic Ruijsenaars-Toda chain [2], see also [1, 14, 15].
It was recently shown in [10] that the elliptic Toda chain is a particular case of the Ruijsenaars chain introduced previously in [16]. The classical $r$-matrix structure for the Ruijsenaars chain was derived in [9] and then adopted to the elliptic Toda chain in [10]. Another important observation of [10] and [16] is that the elliptic Toda chain is gauge equivalent to the discrete Landau-Lifshitz model (or the classical XYZ chain) with some special restrictions. In the next two sections we briefly recall the main results on this topic.
The purpose of the paper is to derive an open elliptic Toda chain with boundary terms using the gauge equivalence mentioned above. Our construction is based on the standard description of the classical transfer-matrix for XYZ chain with boundary [13]. Using the factorized form for the Lax matrices we represent the transfer-matrix as the one describing the open Toda chain with the gauge transformed $K$-matrices. This allows to compute the Hamiltonian, and the integrability follows from integrability of the XYZ chain with boundaries.
2 Lax matrices in factorized form
Introduce the following matrix:
$\begin{array}[]{c}\displaystyle{g(z,{\bf{q}}_{a})=\frac{1}{\theta_{1}(2{\bf{q}}_{a}|\tau)}\left(\begin{array}[]{cc}{\theta_{3}(z-2{\bf{q}}_{a}|2\tau)}&{-\theta_{3}(z+2{\bf{q}}_{a}|2\tau)}\\ \\ {-\theta_{2}(z-2{\bf{q}}_{a}|2\tau)}&{\theta_{2}(z+2{\bf{q}}_{a}|2\tau)}\end{array}\right)\,.}\end{array}$ (2)
This is an intertwining matrix entering the IRF-Vertex correspondence [3] in quantum statistical models. We use the standard notations for Jacobi theta functions. From here on it is assumed that all elliptic functions depend on $\tau$ unless otherwise stated explicitly (like in (2)). Let us introduce a notation for a set of $2\times 2$ matrices:
$\begin{array}[]{c}\displaystyle{L^{[k,m]}(z)=-g^{-1}(z,{\bf{q}}_{k})g(z,{\bf{q}}_{m})e^{P^{m}/2c}\,,\qquad P^{m}={\rm diag}({\bf{p}}_{m},-{\bf{p}}_{m})}\end{array}$ (3)
Direct calculation yields expression for $L^{[k,m]}(z)$ in terms of the elliptic Kronecker function $\phi(z,x)$:
$\begin{array}[]{c}\displaystyle{L^{[k,m]}(z)=\left(\begin{array}[]{cc}{\phi(z,{\bf{q}}_{k}-{\bf{q}}_{m})b^{[k,m]}_{1}}&{\phi(z,{\bf{q}}_{k}+{\bf{q}}_{m})b^{[k,m]}_{2}}\\ {\phi(z,-{\bf{q}}_{k}-{\bf{q}}_{m})b^{[k,m]}_{1}}&{\phi(z,-{\bf{q}}_{k}+{\bf{q}}_{m})b^{[k,m]}_{2}}\end{array}\right)\,,\quad\phi(z,u)=\frac{\theta_{1}^{\prime}(0)\theta_{1}(z+u)}{\theta_{1}(z)\theta_{1}(u)}}\end{array}$ (4)
where
$\begin{array}[]{c}\displaystyle{b^{[k,m]}_{1}=\frac{\theta_{1}({\bf{q}}_{m}-{\bf{q}}_{k})\theta_{1}({\bf{q}}_{m}+{\bf{q}}_{k})}{\theta_{1}^{\prime}(0)\theta_{1}(2{\bf{q}}_{m})}\,e^{{{\bf{p}}_{m}}/{2c}}\ \ b^{[k,m]}_{2}=-\frac{\theta_{1}({\bf{q}}_{m}-{\bf{q}}_{k})\theta_{1}({\bf{q}}_{m}+{\bf{q}}_{k})}{\theta_{1}^{\prime}(0)\theta_{1}(2{\bf{q}}_{m})}\,e^{-{{\bf{p}}_{m}}/{2c}}\,.}\end{array}$ (5)
The Lax matrices for the elliptic Toda chain can be written in several different ways. First possibility is to define the monodromy matrix $T^{\hbox{\tiny{eToda}}}(z)$ as
$\begin{array}[]{c}\displaystyle{T^{\hbox{\tiny{eToda}}}(z)=L^{[n,1]}(z)L^{[1,2]}(z)L^{[2,3]}(z)\dots L^{[n-1,n]}(z)\,,}\end{array}$ (6)
where we use the Lax matrices (4). Another possibility is to use the modified Lax representation [16, 10]:
$\begin{array}[]{c}\displaystyle{{\bf T}^{\hbox{\tiny{eToda}}}(z)={\bf L}^{1}(z){\bf L}^{2}(z)\dots\bf L^{n}(z)\,,\qquad{\bf L}^{a}(z)=\frac{1}{b^{[a-1,a]}_{1}+b^{[a-1,a]}_{2}}\,L^{[a-1,a]}(z)\,.}\end{array}$ (7)
Then ${\bf T}^{\hbox{\tiny{eToda}}}(z)={T}^{\hbox{\tiny{eToda}}}(z)e^{-H^{\hbox{\tiny{eToda}}}/2}$. Explicitly,
$\begin{array}[]{c}\displaystyle{{\bf L}^{a}(z)=\frac{1}{2}\left(\begin{array}[]{cc}{\phi(z,{\bf{q}}_{a-1}-{\bf{q}}_{a})\displaystyle{\frac{e^{{\bf{p}}_{a}/2c}}{\sinh({\bf{p}}_{a}/2c)}}}&{-\phi(z,{\bf{q}}_{a-1}+{\bf{q}}_{a})\displaystyle{\frac{e^{-{\bf{p}}_{a}/2c}}{\sinh({\bf{p}}_{a}/2c)}}}\\ {\phi(z,-{\bf{q}}_{a-1}-{\bf{q}}_{a})\displaystyle{\frac{e^{{\bf{p}}_{a}/2c}}{\sinh({\bf{p}}_{a}/2c)}}}&{-\phi(z,-{\bf{q}}_{a-1}+{\bf{q}}_{a})\displaystyle{\frac{e^{-{\bf{p}}_{a}/2c}}{\sinh({\bf{p}}_{a}/2c)}}}\end{array}\right)\,.}\end{array}$ (8)
Using the identity $\phi(z,u)\phi(z,-u)=\wp(z)-\wp(u)$ we have
$\begin{array}[]{c}\displaystyle{\det{\bf L}^{a}(z)=\frac{1}{4}\frac{1}{\sinh^{2}({\bf{p}}_{a}/2c)}\Big(\wp({\bf{q}}_{a-1}-{\bf{q}}_{a})-\wp({\bf{q}}_{a-1}+{\bf{q}}_{a})\Big)\,.}\end{array}$ (9)
Then it is easy to see that $\log\det{\bf T}^{\hbox{\tiny{eToda}}}(z)$ provides the Hamiltonian (1).
3 Gauge equivalence with XYZ chain
Plugging the factorized representation (3) into (6) gives
$\begin{array}[]{c}\displaystyle{T^{\hbox{\tiny{eToda}}}(z)=(-1)^{n}g^{-1}(z,{\bf{q}}_{n})g(z,{\bf{q}}_{1})e^{P^{1}/2c}g^{-1}(z,{\bf{q}}_{1})g(z,{\bf{q}}_{2})e^{P^{2}/2c}...g^{-1}(z,{\bf{q}}_{n-1})g(z,{\bf{q}}_{n})e^{P^{n}/2c}}\end{array}$ (10)
Therefore, the gauge transformed monodromy matrix ${\mathcal{T}}(z)=g(z,{\bf{q}}_{n})T^{\hbox{\tiny{eToda}}}(z)g^{-1}(z,{\bf{q}}_{n})$ takes the form
$\begin{array}[]{c}\displaystyle{{\mathcal{T}}(z)=(-1)^{n}{\mathcal{L}}^{1}(z){\mathcal{L}}^{2}(z)\dots\mathcal{L}^{n}(z)\,,\qquad{\mathcal{L}}^{k}(z)=g(z,{\bf{q}}_{k})e^{P^{k}/2c}g^{-1}(z,{\bf{q}}_{k})}\end{array}$ (11)
It was explained in [10] that ${\mathcal{L}}^{a}(z)$ are the Lax matrices of the XYZ chain, and ${\mathcal{T}}(z)$ is its standard monodromy matrix [12, 4]. Namely,
$\begin{array}[]{c}\displaystyle{{\mathcal{L}}^{a}(z)=\sigma_{0}{\bf S}^{a}_{0}+\sum\limits_{k=1}^{3}\sigma_{k}\varphi_{k}(z){\bf S}^{a}_{k}\,,\qquad\varphi_{k}(z)=\frac{\theta_{1}^{\prime}(0)\theta_{5-k}(z)}{\theta_{1}(z)\theta_{5-k}(0)}\,,\quad k=1,2,3\,,}\end{array}$ (12)
where $\sigma_{a}$ are the Pauli matrices ($\sigma_{0}$ is $2\times 2$ identity matrix). The variables ${\bf S}^{a}_{k}$ are expressed through ${\bf{p}}_{a}$ and ${\bf{q}}_{a}$ as
$\begin{array}[]{c}\displaystyle{{\bf S}_{0}^{a}=\frac{1}{2}\Big(e^{{\bf{p}}_{a}/2c}+e^{-{\bf{p}}_{a}/2c}\Big)\,,\quad{\bf S}_{k}^{a}=\frac{1}{2}\frac{\theta_{5-k}(0)}{\theta_{1}^{\prime}(0)}\Big(\frac{\theta_{5-k}(2{\bf{q}}_{a})}{\theta_{1}(2{\bf{q}}_{a})}\,e^{{\bf{p}}_{a}/2c}-\frac{\theta_{5-k}(2{\bf{q}}_{a})}{\theta_{1}(2{\bf{q}}_{a})}\,e^{-{\bf{p}}_{a}/2c}\Big)}\end{array}$ (13)
for all a=1,,na=1,...,n and $k=1,2,3$. The Poisson brackets between ${\bf S}_{k}^{a}$ computed through $\{{\bf{p}}_{a},{\bf{q}}_{b}\}=\delta_{ab}$ reproduce the classical Sklyanin algebra [11]
$\begin{array}[]{c}\displaystyle{c\{{\bf S}^{a}_{i},{\bf S}^{a}_{j}\}=-\imath\varepsilon_{ijk}{\bf S}^{a}_{0}{\bf S}^{a}_{k}\,,}\\ \\ \displaystyle{c\{{\bf S}^{a}_{0},{\bf S}^{a}_{i}\}=-\imath\varepsilon_{ijk}{\bf S}^{a}_{j}{\bf S}^{a}_{k}\big(\wp(\omega_{j})-\wp(\omega_{k})\big)\,,}\end{array}$ (14)
where the half-periods $\omega_{j}$
$\begin{array}[]{c}\displaystyle{\omega_{1}=\frac{\tau}{2}\,,\quad\omega_{2}=\frac{1+\tau}{2}\,,\quad\omega_{3}=\frac{1}{2}}\end{array}$ (15)
are numerated according to numeration of the Pauli matrices. The Sklyanin algebra (14) is generated by the classical quadratic exchange relation
$\begin{array}[]{c}\displaystyle{\{{\mathcal{L}}_{1}^{a}(z),{\mathcal{L}}_{2}^{b}(w)\}=\frac{\delta^{ab}}{c}\,[{\mathcal{L}}_{1}^{a}(z){\mathcal{L}}_{2}^{a}(w),r_{12}(z-w)]\,,\qquad r_{12}(z-w)=\frac{1}{2}\sum\limits_{k=1}^{3}\varphi_{k}(z-w)\sigma_{k}\otimes\sigma_{k}\,.}\end{array}$ (16)
The representation (13) for the variables ${\bf S}^{a}_{k}$ corresponds to special values of the Casimir functions ${\bf C}_{1}^{a}=({\bf S}^{a}_{1})^{2}+({\bf S}^{a}_{2})^{2}+({\bf S}^{a}_{3})^{2}$ and ${\bf C}_{2}^{a}=({\bf S}^{a}_{0})^{2}+\sum\limits_{k=1}^{3}({\bf S}^{a}_{k})^{2}\wp(\omega_{k})$:
$\begin{array}[]{c}\displaystyle{{\bf C}_{1}^{a}=0\,,\qquad{\bf C}_{2}^{a}=1\,.}\end{array}$ (17)
4 XYZ chain with boundaries
Following [13] consider the monodromy matrix:
$\begin{array}[]{c}\displaystyle{{\mathcal{T}}^{\hbox{\tiny{openXYZ}}}(z)=K^{+}(z){\mathcal{T}}(z)K^{-}(z){\mathcal{T}}^{-1}(-z)\,,}\end{array}$ (18)
where ${\mathcal{T}}(z)$ (11) satisfies the quadratic classical exchange relation
$\begin{array}[]{c}\displaystyle{c\{{\mathcal{T}}_{1}(z),{\mathcal{T}}_{2}(w)\}=[{\mathcal{T}}_{1}(z){\mathcal{T}}_{2}(w),r_{12}(z-w)]}\end{array}$ (19)
due to (16), and $K^{\pm}(z)$ are $2\times 2$ matrices solving the classical quadratic reflection equation
${[K^{\pm}_{1}(z)K^{\pm}_{2}(w),r_{12}(z-w)]+\,K^{\pm}_{2}(w)r_{12}(z+w)K^{\pm}_{1}(z)-\,K^{\pm}_{1}(z)r_{12}(z+w)K^{\pm}_{2}(w)=0}$ (20)
with the elliptic $r$-matrix from (16). Then the classical transfer-matrix ${\rm tr}{\mathcal{T}}^{\hbox{\tiny{openXYZ}}}(z)$ is a generating function of commuting Hamiltonians since
$\begin{array}[]{c}\displaystyle{\{{\rm tr}{\mathcal{T}}^{\hbox{\tiny{openXYZ}}}(z),{\rm tr}{\mathcal{T}}^{\hbox{\tiny{openXYZ}}}(w)\}=0\,.}\end{array}$ (21)
Notice that $({{\mathcal{L}}}^{a})^{-1}(-z)={{\mathcal{L}}}^{a}(z)/\det({{\mathcal{L}}}^{a}(z))$ since the functions $\varphi_{k}(z)$ (12) are odd in $z$. On the other hand, $\det{{\mathcal{L}}}^{a}(z)={\bf C}_{2}^{a}-\wp(z){\bf C}_{1}^{a}=1$ due to (17). Therefore, ${\mathcal{T}}^{-1}(-z)=(-1)^{n}{\mathcal{L}}^{n}(z){\mathcal{L}}^{n-1}(z)\dots\mathcal{L}^{1}(z)\,.$ Finally,
$\begin{array}[]{c}\displaystyle{{\mathcal{T}}^{\hbox{\tiny{openXYZ}}}(z)=K^{+}(z){\mathcal{L}}^{1}(z){\mathcal{L}}^{2}(z)\dots\mathcal{L}^{n}(z)K^{-}(z){\mathcal{L}}^{n}(z){\mathcal{L}}^{n-1}(z)\dots\mathcal{L}^{1}(z)\,.}\end{array}$ (22)
5 Gauge transformed $K$-matrices
Plugging the factorized representation ${\mathcal{L}}^{k}(z)=g(z,{\bf{q}}_{k})e^{P^{k}/2c}g^{-1}(z,{\bf{q}}_{k})$ (11) into (22) we get
$\begin{array}[]{c}\displaystyle{{\mathcal{T}}^{\hbox{\tiny{openXYZ}}}(z)=K^{+}(z)g(z,{\bf{q}}_{1})e^{P^{1}/2c}g^{-1}(z,{\bf{q}}_{1})\dots g(z,{\bf{q}}_{n})e^{P^{n}/2c}g^{-1}(z,{\bf{q}}_{n})\times}\\ \\ \displaystyle{\times K^{-}(z)g(z,{\bf{q}}_{n})e^{P^{n}/2c}g^{-1}(z,{\bf{q}}_{n})\dots g(z,{\bf{q}}_{1})e^{P^{1}/2c}g^{-1}(z,{\bf{q}}_{1})\,.}\end{array}$ (23)
Consider the gauge transformed monodromy matrix ${\mathcal{T}}^{\hbox{\tiny{openToda}}}(z)=g^{-1}(z,{\bf{q}}_{1}){\mathcal{T}}^{\hbox{\tiny{openXYZ}}}(z)g(z,{\bf{q}}_{1})$:
$\begin{array}[]{c}\displaystyle{{\mathcal{T}}^{\hbox{\tiny{openToda}}}(z)={\tilde{K}}^{+}(z)e^{P^{1}/2c}L^{[1,2]}(z)\dots L^{[n-1,n]}(z){\tilde{K}}^{-}(z)e^{P^{n}/2c}L^{[n,n-1]}(z)\dots L^{[2,1]}(z)\,,}\end{array}$ (24)
where the notation (3) is used and
$\begin{array}[]{c}\displaystyle{{\tilde{K}}^{+}(z)=g^{-1}(z,{\bf{q}}_{1}){K}^{+}(z)g(z,{\bf{q}}_{1})\,,\qquad{\tilde{K}}^{-}(z)=g^{-1}(z,{\bf{q}}_{n}){K}^{-}(z)g(z,{\bf{q}}_{n})\,.}\end{array}$ (25)
The elliptic $K$-matrices solving the reflection equation (20) are known [5]. We write them in the following form:
$\begin{array}[]{c}\displaystyle{K^{\pm}(z)=\sigma_{0}\nu_{0}+\sum\limits_{\alpha=1}^{3}\frac{\nu^{\pm}_{\alpha}}{c_{\alpha}(\tau)}\,\frac{1}{\varphi_{\alpha}(z)}\,\sigma_{\alpha}\,,\qquad c_{k}(\tau)=\Big(\frac{\theta_{5-k}(0)}{\theta_{1}^{\prime}(0)}\Big)^{2}\,,\quad k=1,2,3\,.}\end{array}$ (26)
Equivalently,
$\begin{array}[]{c}\displaystyle{K^{\pm}(z)=\sigma_{0}\nu^{\pm}_{0}+\sigma_{1}\nu^{\pm}_{1}\varphi_{1}(z+\omega_{1})-\sigma_{2}\nu^{\pm}_{2}\varphi_{2}(z+\omega_{2})-\sigma_{3}\nu^{\pm}_{3}\varphi_{3}(z+\omega_{3})\,.}\end{array}$ (27)
In order to compute the gauged transformed $K$-matrices (25) one should use identities with doubled nodular parameter $2\tau$:
$\begin{array}[]{c}\displaystyle{\theta_{2}(x+y|2\tau)\theta_{2}(x-y|2\tau)=\frac{1}{2}\Big(\theta_{3}(x|\tau)\theta_{3}(y|\tau)-\theta_{4}(x|\tau)\theta_{4}(y|\tau)\Big)\,,}\\ \\ \displaystyle{\theta_{2}(x+y|2\tau)\theta_{3}(x-y|2\tau)=\frac{1}{2}\Big(\theta_{2}(x|\tau)\theta_{2}(y|\tau)-\theta_{1}(x|\tau)\theta_{1}(y|\tau)\Big)\,,}\\ \\ \displaystyle{\theta_{3}(x+y|2\tau)\theta_{3}(x-y|2\tau)=\frac{1}{2}\Big(\theta_{3}(x|\tau)\theta_{3}(y|\tau)+\theta_{4}(x|\tau)\theta_{4}(y|\tau)\Big)\,.}\end{array}$ (28)
In particular, it provides $\det g(z,{\bf{q}}_{1})=-\theta_{1}(z)\theta_{1}(2{\bf{q}}_{1})$. Direct computation yields
$\begin{array}[]{c}\displaystyle{{\tilde{K}}^{+}(z)=\left(\begin{array}[]{cc}{\nu^{+}_{0}}&{0}\\ {0}&{\nu^{+}_{0}}\end{array}\right)+\sum\limits_{k=1}^{3}\nu^{+}_{k}\left(\begin{array}[]{cc}{-\varphi_{k}(2{\bf{q}}_{1})}&{\varphi_{k}(2{\bf{q}}_{1},z+\omega_{k})}\\ {\varphi_{k}(-2{\bf{q}}_{1},z+\omega_{k})}&{\varphi_{k}(2{\bf{q}}_{1})}\end{array}\right)\,,}\end{array}$ (29)
where $\varphi_{k}(z,x+\omega_{k})=\frac{\theta_{1}^{\prime}(0)\theta_{5-k}(z+x)}{\theta_{1}(z)\theta_{5-k}(x)}$ for $k=1,2,3$, that is $\varphi_{k}(z,x+\omega_{k})|_{x=0}=\varphi_{k}(z)$. The expression for ${\tilde{K}}^{-}(z)$ is the same as (29) with $\nu^{+}_{k}$ being replaced with $\nu^{-}_{k}$ and ${\bf{q}}_{1}$ being replaced with ${\bf{q}}_{n}$.
6 Open elliptic Toda chain with boundaries
By construction,
$\begin{array}[]{c}\displaystyle{\{{\rm tr}{\mathcal{T}}^{\hbox{\tiny{openToda}}}(z),{\rm tr}{\mathcal{T}}^{\hbox{\tiny{openToda}}}(w)\}=\{{\rm tr}{\mathcal{T}}^{\hbox{\tiny{openXYZ}}}(z),{\rm tr}{\mathcal{T}}^{\hbox{\tiny{openXYZ}}}(w)\}=0\,.}\end{array}$ (30)
Let us compute the Hamiltonian for the open elliptic Toda chain generated by ${\rm tr}{\mathcal{T}}^{\hbox{\tiny{openToda}}}(z)$ (24). Namely, consider the coefficient behind $1/z^{2n-2}$ in the expansion of ${\rm tr}{\mathcal{T}}^{\hbox{\tiny{openToda}}}(z)$ near $z=0$:
$\begin{array}[]{c}\displaystyle{H=\mathop{\hbox{Res}}\limits\limits_{z=0}z^{2n-1}{\rm tr}{\mathcal{T}}^{\hbox{\tiny{openToda}}}(z)\,.}\end{array}$ (31)
Notice that matrices $L^{[a,b]}(z)$ (4) have simple poles at $z=0$, while the $K$-matrices ${\tilde{K}}^{\pm}(z)$ (29) are regular at $z=0$. Therefore,
$\begin{array}[]{c}\displaystyle{H={\rm tr}\Big({\tilde{K}}^{+}(0)e^{P^{1}/2c}A^{[1,2]}\dots A^{[n-1,n]}{\tilde{K}}^{-}(0)e^{P^{n}/2c}A^{[n,n-1]}\dots A^{[2,1]}\Big)\,,\quad A^{[k,m]}=\mathop{\hbox{Res}}\limits\limits_{z=0}L^{[k,m]}(z)\,.}\end{array}$ (32)
The residues $A^{[k,m]}$ are easily calculated using the property $\mathop{\hbox{Res}}\limits\limits_{z=0}\phi(z,u)=1$:
$\begin{array}[]{c}\displaystyle{A^{[k,m]}=\left(\begin{array}[]{c}1\\ 1\end{array}\right)\otimes(b_{1}^{[k,m]}\ b_{2}^{[k,m]})\,.}\end{array}$ (33)
It is also useful to mention that
$\begin{array}[]{c}\displaystyle{{\tilde{K}}^{+}(0)=\left(\begin{array}[]{cc}{\nu^{+}_{0}}&{0}\\ {0}&{\nu^{+}_{0}}\end{array}\right)-{\tilde{f}}^{+}({\bf{q}}_{1})\left(\begin{array}[]{c}1\\ 1\end{array}\right)\otimes(1\ -1)\,,\quad{\tilde{f}}^{+}({\bf{q}}_{1})=\sum\limits_{k=1}^{3}\nu^{+}_{k}\varphi_{k}(2{\bf{q}}_{1})}\end{array}$ (34)
and similarly
$\begin{array}[]{c}\displaystyle{{\tilde{K}}^{-}(0)=\left(\begin{array}[]{cc}{\nu^{+}_{0}}&{0}\\ {0}&{\nu^{-}_{0}}\end{array}\right)-{\tilde{f}}^{-}({\bf{q}}_{n})\left(\begin{array}[]{c}1\\ 1\end{array}\right)\otimes(1\ -1)\,,\quad{\tilde{f}}^{-}({\bf{q}}_{n})=\sum\limits_{k=1}^{3}\nu^{-}_{k}\varphi_{k}(2{\bf{q}}_{n})\,.}\end{array}$ (35)
In particular, this yields
$\begin{array}[]{c}\displaystyle{(b_{1}^{[2,1]}\ b_{2}^{[2,1]}){\tilde{K}}^{+}(0)e^{P^{1}/2c}\left(\begin{array}[]{c}1\\ 1\end{array}\right)=2\frac{\theta_{1}({\bf{q}}_{1}-{\bf{q}}_{2})\theta_{1}({\bf{q}}_{1}+{\bf{q}}_{2})}{\theta_{1}^{\prime}(0)\theta_{1}(2{\bf{q}}_{1})}\sinh\big(\frac{{\bf{p}}_{1}}{2c}\big)\Big(2\nu_{0}^{+}\cosh\big(\frac{{\bf{p}}_{1}}{2c}\big)-{\tilde{f}}^{+}({\bf{q}}_{1})\Big)}\end{array}$ (36)
One may use an alternatively expression for ${\tilde{f}}^{+}({\bf{q}}_{1})$ (and similarly for ${\tilde{f}}^{-}({\bf{q}}_{n})$) coming from identity
$\begin{array}[]{c}\displaystyle{\big({\tilde{f}}^{+}({\bf{q}}_{1})\big)^{2}=\sum\limits_{k=0}^{3}({\bar{\nu}}_{k}^{+})^{2}\wp({\bf{q}}_{1}+\omega_{k})-\sum\limits_{k=1}^{3}({\nu}_{k}^{+})^{2}\wp(\omega_{k})\,,}\end{array}$ (37)
where
[…]
Finally, we come to the following Hamiltonian:
$\begin{array}[]{c}\displaystyle{H^{\hbox{\tiny{open}}}=\log\frac{1}{\sinh({\bf{p}}_{1}/2c)}-\log\Big(2\nu_{0}^{+}\cosh({\bf{p}}_{1}/2c)-{\tilde{f}}^{+}({\bf{q}}_{1})\Big)+}\\ \\ \displaystyle{+\log\frac{1}{\sinh({\bf{p}}_{n}/2c)}-\log\Big(2\nu_{0}^{-}\cosh({\bf{p}}_{n}/2c)-{\tilde{f}}^{-}({\bf{q}}_{n})\Big)+}\\ \\ \displaystyle{+\sum\limits_{a=2}^{n-1}\log\frac{1}{\sinh^{2}({\bf{p}}_{a}/2c)}+\sum\limits_{a=2}^{n}\log\Big(\wp({\bf{q}}_{a-1}-{\bf{q}}_{a})-\wp({\bf{q}}_{a-1}+{\bf{q}}_{a})\Big)\,.}\end{array}$ (39)
Consider two particular cases.
Example 1: pure open chain. Let $\nu^{\pm}_{k}=0$, $k=1,2,3$ and $\nu^{\pm}_{0}=1$. Then we come to a ”pure” open elliptic Toda chain (corresponding to $K^{\pm}(z)=\sigma_{0}$ in XYZ chain):
$\begin{array}[]{c}\displaystyle{H^{\hbox{\tiny{pure\ open}}}=\log\frac{1}{\sinh({\bf{p}}_{1}/2c)\cosh({\bf{p}}_{1}/2c)}+\log\frac{1}{\sinh({\bf{p}}_{n}/2c)\cosh({\bf{p}}_{n}/2c)}+}\\ \displaystyle{+\sum\limits_{a=2}^{n-1}\log\frac{1}{\sinh^{2}({\bf{p}}_{a}/2c)}+\sum\limits_{a=2}^{n}\log\Big(\wp({\bf{q}}_{a-1}-{\bf{q}}_{a})-\wp({\bf{q}}_{a-1}+{\bf{q}}_{a})\Big)\,.}\end{array}$ (40)
Example 2: boundary terms. Let $\nu^{\pm}_{0}=0$. Then
$\begin{array}[]{c}\displaystyle{H^{\hbox{\tiny{boundary}}}=\log\frac{1}{\sinh({\bf{p}}_{1}/2c)}-\frac{1}{2}\log\Big(\sum\limits_{k=0}^{3}({\bar{\nu}}_{k}^{+})^{2}\wp(q_{1}+\omega_{k})-\sum\limits_{k=1}^{3}({\nu}_{k}^{+})^{2}\wp(\omega_{k})\Big)+}\\ \displaystyle{+\log\frac{1}{\sinh({\bf{p}}_{n}/2c)}-\frac{1}{2}\log\Big(\sum\limits_{k=0}^{3}({\bar{\nu}}_{k}^{-})^{2}\wp(q_{n}+\omega_{k})-\sum\limits_{k=1}^{3}({\nu}_{k}^{-})^{2}\wp(\omega_{k})\Big)+}\\ \displaystyle{+\sum\limits_{a=2}^{n-1}\log\frac{1}{\sinh^{2}({\bf{p}}_{a}/2c)}+\sum\limits_{a=2}^{n}\log\Big(\wp({\bf{q}}_{a-1}-{\bf{q}}_{a})-\wp({\bf{q}}_{a-1}+{\bf{q}}_{a})\Big)\,.}\end{array}$ (41)
In conclusion, we note again that the Ruijsenaars-Toda chain [2] generalizes the elliptic Toda chain (1), and it is gauge equivalent to a generic XYZ chain (without restrictions (17)). Similarly to the results of the present paper it is then possible to derive the open Ruijsenaars-Toda chain with boundaries. The boundary $K$-matrices can be also chosen to be dynamical, see the Zhukovsky-Volterra gyrostat [7] and the 4-constant ${\rm BC}_{1}$ Ruijsenaars-van Diejen model [8]. These generalizations will be studied elsewhere.
Acknowledgments.
This work was supported by the Russian Science Foundation under grant no. 25-11-00081, https://rscf.ru/project/25-11-00081/ .